# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

### v1

@@ 5 Result
We evaluate the performance of standard PPO and AdaRFT (PPO) across multiple training setups and two model sizes: Qwen 2.5 MATH 1.5B and Qwen 2.5 7B. Figure 3 presents the learning curves averaged across six benchmarks, while Table 1 and 2 provide a detailed breakdown of accuracy and training efficiency at step 60 and step 100. On average, models trained with AdaRFT (PPO) outperform their PPO-only counterparts in both final accuracy and training efficiency. This improvement is particularly notable in non-uniform data distributions, where curriculum adaptation is most beneficial.
@@ 5.1 Training Efficiency
As shown in Figure 3 and Table 1, models with AdaRFT (PPO) require substantially fewer training steps to reach the same level of performance as standard PPO. Specifically, we measure how many additional training steps PPO requires to match the performance that AdaRFT (PPO) attains at step 60. Because evaluation is performed only every 5 steps during training, we estimate these step savings by interpolating the smoothed accuracy curves. We apply exponential smoothing with a factor of 0.7 to reduce evaluation noise and obtain stable estimates of performance over time.
For Qwen 2.5 Math 1.5B, PPO requires 26 additional steps in the skew-difficult setting and 29 in the uniform setting to reach the performance achieved by AdaRFT (PPO) at step 60—corresponding to 43.7% and 48.3% reductions in training steps, respectively. Even in the skew-easy setting, PPO lags behind by 9 steps. The efficiency gains are even more pronounced with the larger Qwen 2.5 7B model: in the skew-difficult setting, PPO requires 44 additional steps, a 73.2% increase, and in the skew-easy setting, 64 additional steps, a 106.5% increase. On uniformly sampled data, both methods converge at a similar rate, though AdaRFT (PPO) maintains a slight advantage.
Notably, these gains are accompanied by a reduction in computational cost per step. As shown in Table 1, AdaRFT (PPO) consistently achieves lower average time per training step than standard PPO across all settings. While the underlying causes are discussed in Section 6, this result highlights that AdaRFT improves not only convergence speed but also overall training throughput, making it both sample- and compute-efficient.
| Model | Setup | Method |
| Qwen 2.5 Math 1.5B | skew-difficult | PPO | 69.14 | 60.20 | 20.36 | 12.13 | 45.31 | 7.50 | 35.77 | 26 (43.67%) | 134.44 | ||||||||||||||||
| AdaRFT | 71.42 | 63.00 | 20.51 | 15.07 | 45.62 | 10.83 | 37.74 | 124.66 | |||||||||||||||||||
| uniform | PPO | 70.74 | 61.20 | 19.61 | 14.34 | 45.00 | 11.67 | 37.09 | 29 (48.32%) | 129.24 | |||||||||||||||||
| AdaRFT | 72.55 | 65.20 | 21.10 | 13.24 | 47.19 | 14.17 | 38.91 | 120.21 | |||||||||||||||||||
| skew-easy | PPO | 70.58 | 61.40 | 20.95 | 14.34 | 47.50 | 12.92 | 37.95 | 9 (14.62%) | 119.99 | |||||||||||||||||
| AdaRFT | 72.25 | 63.60 | 19.76 | 15.81 | 46.88 | 14.17 | 38.74 | 120.91 | |||||||||||||||||||
| Qwen 2.5 7B | skew-difficult | PPO | 88.17 | 69.40 | 21.99 | 25.00 | 45.63 | 9.58 | 43.29 | 44 (73.18%) | 258.87 | ||||||||||||||||
| AdaRFT | 89.08 | 72.20 | 24.52 | 24.26 | 48.12 | 13.33 | 45.25 | 247.73 | |||||||||||||||||||
| uniform | PPO | 89.61 | 71.00 | 23.92 | 23.53 | 49.06 | 9.17 | 44.38 | 3 (5.27%) | 254.38 | |||||||||||||||||
| AdaRFT | 88.32 | 73.40 | 24.81 | 22.06 | 45.63 | 15.83 | 45.01 | 241.64 | |||||||||||||||||||
| skew-easy | PPO | 88.86 | 72.80 | 24.07 | 22.79 | 48.13 | 13.33 | 45.00 | 64 (106.5%) | 250.22 | |||||||||||||||||
| AdaRFT | 88.48 | 73.40 | 23.77 | 23.53 | 49.69 | 12.92 | 45.30 | 253.21 |
@@ 5.2 Model Performance
In addition to improving efficiency, AdaRFT (PPO) also improves the final model performance. As shown in Table 2, at the end of training (step 100), curriculum learning yields consistent improvements in final accuracy across all configurations. As shown in Table 2, AdaRFT (PPO) outperforms standard PPO in nearly every setting, with especially strong gains on the Qwen 2.5 Math 1.5B model. The reported averages reflect accuracy across six diverse benchmarks: GSM8K, MATH 500, OlympiadBench, Minerva Math, AMC 23, and AIME 24.
@@
On skew-difficult data, the Qwen 2.5 Math 1.5B model improves from 37.41% (PPO) to 40.48% (AdaRFT (PPO)), a gain of over 3 percentage points in average accuracy. Similar improvements appear in the uniform setting, where AdaRFT (PPO) reaches 41.11%, compared to 37.20% with PPO. Even on skew-easy data, where the baseline performs well, curriculum learning still improves performance, reaching 39.18% versus 38.46%. For the larger Qwen 2.5 7B model, final accuracy gains are also consistent, though slightly more modest. In the skew-difficult setting, AdaRFT (PPO) improves from 44.17% to 46.83%. In the uniform setting, accuracy rises from 44.70% to 46.92%, and in the skew-easy case, from 45.07% to 45.94%. These results show that curriculum learning is effective even for stronger models, enhancing both stability and peak performance.
@@ 6.1 Analysis on Efficiency and Accuracy
AdaRFT also results in slightly lower average training time per step. This is largely due to the fact that easier problems require fewer tokens to solve. For example, an arithmetic reasoning question from GSM8K might require only around 200 tokens for Qwen 2.5 7B to reach a correct answer, whereas a competition-level math problem from AIME could require around 2000 tokens—a 10 difference in rollout length. The total token length affects multiple components of the training step, including the rollout itself and the subsequent PPO update. While PPO update time does not scale linearly with sequence length due to batching and attention computation patterns, longer sequences still incur higher compute cost. As a result, curriculum learning’s tendency to prioritize shorter, easier problems early in training leads to shorter sequences on average, reducing per-step compute and improving overall training throughput.
@@ 6.2 Data Difficulty on Model Performance
To better understand the effect of data difficulty on model performance, we introduce two additional data distributions: easy-extreme and hard-extreme. Unlike the skew-difficult and skew-easy distributions, which still include a mix of difficulty levels, the easy-extreme and hard-extreme sets consist exclusively of the most polarized examples. Specifically, easy-extreme contains only the easiest samples with difficulty levels no greater than 15, while hard-extreme includes only the hardest samples with difficulty levels of at least 97. Each of these extreme distributions consists of approximately 8,000 samples, providing a focused and controlled evaluation of model behavior under minimal or maximal difficulty conditions. We trained a Qwen 2.5 7B model on each of the two extreme distributions using PPO, and compared their performance to models trained on the uniform distribution with PPO (Uniform) and with AdaRFT instantiated with PPO (Uniform + AdaRFT), as described in Section 5. The results are presented in Figure 4.
@@
The leftmost panel shows that uniform + AdaRFT achieves the highest overall accuracy throughout training, outperforming both uniform and the two extreme settings. This highlights the effectiveness of AdaRFT in guiding the model through an optimal difficulty progression. In contrast, hard-extreme struggles significantly, with a flat and lower trajectory, indicating that exposing the model only to very difficult problems limits learning progress. This suggests that without a gradual exposure strategy, models trained on only the hardest problems are unable to bootstrap their capabilities effectively.
@@ Acknowledgement
Future Works
Despite these encouraging results, there are important limitations. Our curriculum strategy estimates task difficulty by measuring the model’s empirical success rate over 128 attempts per problem, effectively using average success as a proxy for difficulty. While this approach is simple and model-grounded, it is computationally expensive and may not capture the full spectrum of complexity. More sophisticated or efficient alternatives could include model-based difficulty estimation, where a lightweight predictive model learns to estimate difficulty from features such as response length, number of reasoning steps, or uncertainty signals—bypassing the need for repeated sampling. Such approaches could enable faster, more scalable curriculum updates and better generalization across tasks. Future work could explore more adaptive curriculum schedules using richer signals (e.g., loss trends, confidence, or gradient variance), as well as co-adapting the model or optimizer with the curriculum. Finally, understanding when curriculum learning offers diminishing returns, as observed in balanced settings or high-capacity models, is crucial for developing general-purpose and compute-aware curricula.
@@ A.1 Training Configuration
We trained both the actor and critic models using the PPO algorithm. Training was conducted on a single node equipped with 8 A100 GPUs. Each model was trained for approximately 100 steps using the veRL library (Sheng et al., 2024). The key training configuration for Qwen 2.5 7B is shown below. For Qwen 2.5 MATH 1.5B, we set max_response_length=3000 and ulysses_sequence_parallel_size=1 to accommodate its shorter context window of 4096 tokens, while keeping all other parameters unchanged.
export VLLM_ATTENTION_BACKEND=XFORMERS
python3 -m verl.trainer.main_ppo \
data.train_files="$train_files" \
data.val_files="$test_files" \
data.train_batch_size=1024 \
data.max_prompt_length=1024 \
data.max_response_length=8000 \
actor_rollout_ref.model.path=Qwen2.5-7B \
actor_rollout_ref.actor.optim.lr=1e-6 \
actor_rollout_ref.model.use_remove_padding=True \
actor_rollout_ref.actor.ulysses_sequence_parallel_size=2 \
actor_rollout_ref.model.enable_gradient_checkpointing=True \
actor_rollout_ref.actor.ppo_mini_batch_size=1024 \
actor_rollout_ref.actor.use_dynamic_bsz=True \
actor_rollout_ref.actor.ppo_max_token_len_per_gpu=8000 \
actor_rollout_ref.actor.fsdp_config.param_offload=False \
actor_rollout_ref.actor.fsdp_config.optimizer_offload=False \
actor_rollout_ref.rollout.tensor_model_parallel_size=2 \
actor_rollout_ref.rollout.name=vllm \
actor_rollout_ref.rollout.gpu_memory_utilization=0.5 \
actor_rollout_ref.rollout.log_prob_max_token_len_per_gpu=8000 \
actor_rollout_ref.ref.fsdp_config.param_offload=True \
actor_rollout_ref.ref.log_prob_max_token_len_per_gpu=8000 \
critic.optim.lr=1e-5 \
critic.ulysses_sequence_parallel_size=2 \
critic.model.use_remove_padding=True \
critic.model.path=Qwen2.5-7B \
critic.model.enable_gradient_checkpointing=True \
critic.ppo_max_token_len_per_gpu=30000 \
critic.model.fsdp_config.param_offload=False \
critic.model.fsdp_config.optimizer_offload=False \
algorithm.kl_ctrl.kl_coef=0.001 \
trainer.critic_warmup=0 \
trainer.logger=[’console’,’wandb’] \
trainer.project_name=’verl_examples’ \
trainer.experiment_name=’Qwen2.5-7B’ \
trainer.n_gpus_per_node=8 \
+trainer.val_before_train=True \
trainer.nnodes=1 \
trainer.save_freq=10 \
trainer.test_freq=5 \
trainer.total_epochs=15 $@
@@ A.2 AdaRFT Parameters
To enable effective AdaRFT, we tuned several hyperparameters: the training batch size was set to , the target reward to , the sensitivity parameter to , the step size to , and the initial target difficulty to .
The curriculum update mechanism is governed by the difference between the current average reward and the target reward . The difficulty adjustment is computed using a scaled hyperbolic tangent function:
This formulation yields approximately linear updates when , promoting stable learning. When the reward diverges significantly from the target, the function saturates, effectively limiting the magnitude of difficulty updates and preventing destabilizing jumps.
Because the reward is bounded in and the difficulty metric in , we set the step size to ensure alignment in scale. The modulation parameter provides a smooth and controlled progression of difficulty throughout training.

### v2

@@ 5 Result
We evaluate the performance of standard PPO and AdaRFT (PPO) across multiple training setups and two model sizes: Qwen 2.5 MATH 1.5B and Qwen 2.5 7B. Figure 3 presents the learning curves averaged across six benchmarks, while Table 1 and 2 provide a detailed breakdown of accuracy and training efficiency. On average, models trained with AdaRFT (PPO) outperform their PPO-only counterparts in both final accuracy and training efficiency. This improvement is particularly notable in non-uniform data distributions, where curriculum adaptation is most beneficial.
@@ 5.1 Training Efficiency
As shown in Figure 3 and Table 1, models trained with AdaRFT consistently require fewer training steps to match the performance of those trained with standard PPO or PPO on filtered data. Specifically, we report how many additional steps are needed for PPO variants to match the performance of AdaRFT at step 60 for Qwen 2.5 Math 1.5B, and step 40 for Qwen 2.5 7B. Because models are evaluated only every 5 training steps, we apply exponential smoothing with a smoothing parameter to the accuracy curves to reduce variance and obtain stable estimates of performance over time. The shaded areas in Figure 3 represent the raw, unsmoothed accuracy 1%, offering a visual cue for the typical fluctuation in evaluation accuracy.
Model | Setup | Method | Avg Step Time (s) | Extra Steps (%) |
Qwen2.5 Math 1.5B | skew-difficult | AdaRFT | 122.24 | +0 (0.0%) |
PPO | 132.95 | +43 (71.7%) | ||
PPO (w/ Filter) | 128.20 | +49 (81.7%) | ||
uniform | AdaRFT | 121.31 | +0 (0.0%) | |
PPO | 126.82 | +34 (56.7%) | ||
PPO (w/ Filter) | 126.35 | +52 (86.7%) | ||
skew-easy | AdaRFT | 120.52 | +0 (0.0%) | |
PPO | 121.15 | +16 (26.7%) | ||
PPO (w/ Filter) | 115.12 | +21 (35.0%) | ||
Qwen2.5 7B | skew-difficult | AdaRFT | 239.92 | +0 (0.0%) |
PPO | 246.21 | +24 (60.0%) | ||
PPO (w/ Filter) | 254.22 | +25 (62.5%) | ||
uniform | AdaRFT | 234.16 | +0 (0.0%) | |
PPO | 243.82 | +13 (32.5%) | ||
PPO (w/ Filter) | 263.11 | +23 (57.5%) | ||
skew-easy | AdaRFT | 247.44 | +0 (0.0%) | |
PPO | 235.27 | +20 (50.0%) | ||
PPO (w/ Filter) | 233.13 | +17 (42.5%) |
For Qwen 2.5 Math 1.5B, standard PPO requires 43 extra steps (+71.7%) in the skew-difficult setting and 34 steps (+56.7%) in the uniform setting to match AdaRFT’s performance. PPO with filtered training data requires even more: +49 steps (81.7%) and +52 steps (86.7%) in the respective settings. In the skew-easy scenario, PPO requires +16 steps (26.7%), while PPO with filtered data needs +21 steps (35.0%) to catch up to AdaRFT. The efficiency gains remain significant with the larger Qwen 2.5 7B model. In the skew-difficult setting, PPO and PPO with filtered data require +24 steps (60.0%) and +25 steps (62.5%), respectively.
In addition to improved sample efficiency, AdaRFT also achieves faster average training time per step across nearly all settings, as reported in Table 1. Though PPO with filtered data can sometimes offer marginal gains in per-step time (e.g., skew-easy setups), it still falls behind in convergence speed. While the underlying causes are discussed in Section 6, these results underscore that AdaRFT is both sample-efficient and compute-efficient, delivering faster and more cost-effective training.
@@ 5.2 Model Performance
In addition to improving efficiency, AdaRFT (PPO) also improves the final model performance. As shown in Table 2, at the end of training (step 100), AdaRFT yields consistent improvements in final accuracy across all configurations. The reported averages reflect accuracy across six diverse benchmarks: GSM8K, MATH 500, OlympiadBench, Minerva Math, AMC 23, and AIME 24.
On skew-difficult data, the Qwen 2.5 Math 1.5B model improves from 37.41% (PPO) to 40.48% (AdaRFT (PPO)), a gain of over 3 percentage points in average accuracy. Similar improvements appear in the uniform setting, where AdaRFT (PPO) reaches 41.11%, compared to 37.20% with PPO. Even on skew-easy data, where the baseline performs well, curriculum learning still improves performance, reaching 39.18% versus 38.46%. For the larger Qwen 2.5 7B model, final accuracy gains are also consistent, though slightly more modest. In the skew-difficult setting, AdaRFT (PPO) improves from 44.17% to 46.83%. In the uniform setting, accuracy rises from 44.70% to 46.92%, and in the skew-easy case, from 45.07% to 45.94%. These results show that AdaRFT is effective even for stronger models, enhancing both stability and peak performance.
@@ 6.1 Analysis on Efficiency and Accuracy
AdaRFT also results in a slightly lower average training time per step. This is largely due to the fact that easier problems require fewer tokens to solve. For example, an arithmetic reasoning question from GSM8K might require only around 200 tokens for Qwen 2.5 7B to reach a correct answer, whereas a competition-level math problem from AIME could require around 2000 tokens—a 10 difference in rollout length. The total token length affects multiple components of the training step, including the rollout itself and the subsequent PPO update. While PPO update time does not scale linearly with sequence length due to batching and attention computation patterns, longer sequences still incur higher compute costs. As a result, curriculum learning’s tendency to prioritize shorter, easier problems early in training leads to shorter sequences on average, reducing per-step compute and improving overall training throughput.
@@ 6.2 Data Difficulty on Model Performance
To better understand the effect of data difficulty on model performance, we introduce two additional data distributions: easy-extreme and hard-extreme. Unlike the skew-difficult and skew-easy distributions, which still include a mix of difficulty levels, the easy-extreme and hard-extreme sets consist exclusively of the most polarized examples. Specifically, easy-extreme contains only the easiest samples with difficulty levels no greater than 15, while hard-extreme includes only the hardest samples with difficulty levels of at least 97. Each of these extreme distributions consists of approximately 8,000 samples, providing a focused and controlled evaluation of model behavior under minimal or maximal difficulty conditions. We trained a Qwen 2.5 7B model on each of the two extreme distributions using PPO, and compared their performance to models trained on the uniform distribution with PPO (Uniform) and with AdaRFT instantiated with PPO (Uniform + AdaRFT), as described in Section 5. The results are presented in Figure 4. The key takeaway is that training on only overly easy or hard problems fails to provide useful learning signals, reinforcing the need for AdaRFT to adaptively steer models toward challenges matched to their current ability.
@@
The leftmost panel of Figure 4 shows that uniform + AdaRFT achieves the highest overall accuracy throughout training, outperforming both uniform and the two extreme settings. This highlights the effectiveness of AdaRFT in guiding the model through an optimal difficulty progression. In contrast, hard-extreme struggles significantly, with a flat and lower trajectory, indicating that exposing the model only to very difficult problems limits learning progress. This suggests that without a gradual exposure strategy, models trained on only the hardest problems are unable to bootstrap their capabilities effectively.
@@
6.3 Ablation on Target Reward
To better understand the role of the target reward in ADARFT, we perform an ablation study varying in the target difficulty update rule. Recall that controls the target average reward the model is expected to achieve and implicitly steers the curriculum: lower values prioritize easier problems, while higher values shift the curriculum toward more challenging samples. We train a Qwen 2.5 Math 1.5B model on the uniform data distribution with AdaRFT (PPO) using three different values of : 0.2, 0.5, and 0.8. For comparison, we also include standard PPO without ADARFT (denoted as “w/o ADARFT”) as a baseline.
As shown in Figure 5, the model trained with achieves the highest accuracy throughout training. This supports our theoretical motivation in Section 3.4: maximizing reward variance, which occurs when success rate , provides the strongest learning signal. Models with and underperform likely due to curriculum misalignment: overly focuses on easy problems, while overemphasizes difficult ones, both of which limit the model’s capacity to generalize. The reward and difficulty curves align with the accuracy outcomes discussed above. The configuration maintains a stable reward near 0.5, reflecting balanced difficulty exposure. In contrast, results in overly high reward (i.e., easy samples), while maintains a reward around 0.2 for most of training, indicating the model is repeatedly presented with overly difficult problems. As expected, response length is the shortest for and longest for , consistent with the idea that longer responses correlate with problem complexity.
6.4 Training on LLM-Estimated Difficulty
In addition to rollout-based difficulty estimation, we explore an alternative strategy that uses LLM-judged difficulty levels to guide curriculum construction. As described in Section 4.1, we prompt GPT-4o (gpt-4o-0806) to assign difficulty levels to math problems in the DeepScaleR dataset according to the AoPS rubric. This approach offers a lightweight and scalable alternative to computing pass@ success rates from model rollouts, making it especially attractive in low-resource scenarios.
To assess the effectiveness of this strategy, we train a Qwen 2.5 Math 1.5B model on the skew-difficult distribution using AdaRFT (PPO) with two curriculum schedules: one based on rollout-derived pass@ difficulty, and the other guided by GPT-4o’s difficulty ratings. Since the LLM-judged difficulty is on a scale of 1 to 5 (rather than 0 to 100), we set the step size hyperparameter to align the difficulty adjustment magnitude with the reward signal. All other hyperparameters are kept unchanged. As shown in Figure 6, both curriculum strategies outperform standard PPO without curriculum learning. While rollout-based difficulty estimation yields the strongest gains, the LLM-judged curriculum still provides a noticeable improvement over the baseline.
These results demonstrate that AdaRFT remains effective even when the difficulty signal is derived from heuristic or approximate sources like LLM judgments. Although less precise than empirical pass@ metrics, the LLM-based difficulty still provides enough structure to enable meaningful curriculum adaptation. This makes it a practical fallback when rollout computation is too costly, and suggests that future work could explore hybrid approaches that combine lightweight heuristics with periodic empirical calibration.
6.5 AdaRFT with Diverse RL Algorithms
To evaluate the generality of AdaRFT beyond PPO, we trained the Qwen 2.5 Math 1.5B model on a skew-difficult data distribution using two alternative reinforcement learning algorithms: REINFORCE++ and GRPO (see implementation details in Appendix A). As shown in Figure 7, AdaRFT significantly improves both the convergence speed and final accuracy across these variants. Across both cases, the adaptive curriculum acts orthogonally to the underlying optimization method. These results reinforce the plug-and-play nature of AdaRFT: it consistently enhances sample efficiency and policy robustness across algorithmic choices, making it broadly applicable in diverse reinforcement finetuning pipelines. Notably, this generalization holds without any additional tuning or algorithm-specific modifications, underscoring the practical utility of curriculum-aware training in both lightweight and computation-heavy RFT settings.
@@ A.1 Training Configuration
We trained both the actor and critic models using the PPO algorithm on a single node with 8 A100 GPUs. Each model was trained for approximately 100 optimization steps using the veRL library (Sheng et al., 2024). We used two model variants: Qwen2.5-7B and Qwen2.5-MATH-1.5B. The latter has a shorted context window, so we adjusted the max response length and the sequence parallel size accordingly.
Table 3 summarizes the core hyperparameter settings used across all three algorithms: PPO, GRPO, and REINFORCE++. We highlight both shared defaults and algorithm-specific overrides, including KL treatment modes, rollout settings, and critic configurations.
@@ A.2 AdaRFT Parameters
To enable effective AdaRFT, we incorporated a simple curriculum learning mechanism that dynamically adjusts task difficulty during training. We tuned several key hyperparameters: training batch size , target reward , sensitivity parameter , step size , and initial difficulty . These settings were used uniformly across all experiments listed in Table 3. The difficulty is updated as a function of the discrepancy between the running average reward and the target reward , using the following rule:
This formulation promotes stable learning by providing approximately linear updates when , and saturated updates when rewards deviate significantly, thereby avoiding abrupt difficulty shifts. Because the reward is bounded in and the difficulty metric spans , we set the step size to align their scales. The modulation parameter ensures smooth and controlled progression throughout training.
Category | Parameter | PPO | GRPO | REINFORCE++ |
Algorithm-Specific Settings | ||||
General | Advantage estimator | GAE | GRPO | REINFORCE++ |
Gamma () | 1.0 | — | — | |
Lambda () | 1.0 | — | — | |
Batch size | 1024 | 1024 | 1024 | |
Max prompt length | 1024 | 1024 | 1024 | |
Gradient checkpointing | Enabled | Enabled | Enabled | |
Actor | Learning rate | |||
Mini-batch size | 1024 | 1024 | 1024 | |
Dynamic batch size | Enabled | Enabled | Enabled | |
KL penalty role | Reward | Loss | Loss | |
KL loss type | Fixed | Low-variance KL | MSE | |
KL loss coefficient () | 0.001 | 0.001 | 0.001 | |
Entropy coefficient | 0.001 | 0.001 | 0 | |
Clip ratio | 0.2 | 0.2 | 0.2 | |
Gradient clipping | 1.0 | 1.0 | 1.0 | |
Sequence parallel size | Model-specific | Model-specific | Model-specific | |
Rollout | Backend | vLLM | vLLM | vLLM |
Tensor model parallel size | 2 | 2 | 2 | |
Rollouts per sample | 1 | 4 | 1 | |
Nucleus sampling | 1.0 | 1.0 | 1.0 | |
GPU memory utilization | 0.5 | 0.5 | 0.5 | |
Sampling temperature | 1.0 | 1.0 | 1.0 | |
Critic | Warmup steps | 0 | — | — |
Learning rate | — | — | ||
Sequence parallel size | Model-specific | — | — | |
Model-Specific Overrides (shared across all algorithms) | ||||
Qwen2.5-7B | Max response length | 8000 | 8000 | 8000 |
Sequence parallel size | 2 | 2 | 2 | |
Max token length / GPU | 8000 | 8000 | 8000 | |
Qwen2.5-MATH-1.5B | Max response length | 3000 | 3000 | 3000 |
Sequence parallel size | 1 | 1 | 1 | |
Max token length / GPU | 16000 | 16000 | 16000 | |
AdaRFT Parameters | ||||
Curriculum Learning | Target reward () | 0.5 | 0.5 | 0.5 |
Sensitivity () | 2 | 2 | 2 | |
Step size () | 50 | 50 | 50 | |
Initial difficulty () | 0 | 0 | 0 |
A.3 Prompt for Difficulty Estimation Using LLM as a Judge
The prompt used for difficulty estimation (as described in Section 4.1) is shown in Table 4, Table 5, and Table 6. The descriptions of the difficulty scales and examples are sourced from the AoPS Wiki.555https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings Although GPT-4o is prompted to rate problem difficulty on a scale from 1 to 10, we found that over 95% of the problems fall within the range of 1 to 5. Therefore, we clip the scores and use a revised scale from 1 to 5. In addition to integer scores, we also allow half-point increments such as 1.5, 2.0, and 2.5 for finer-grained difficulty estimation.